OpenStudy (anonymous):

determine the srea of the region bounded by f(x)=8x^2+4 and the x-axis is between 2 and 3

OpenStudy (amistre64):

just integrate f(x) from 2 to 3

OpenStudy (anonymous):

can you show me step by step, im really confused

OpenStudy (amistre64):

as far as im reading it right .... do you know what integration is?

OpenStudy (anonymous):

is it the anti derivative

OpenStudy (amistre64):

correct so, whatis the antiD of 8x^2? and what it is it of 4?

OpenStudy (anonymous):

(8/3)x^3?

OpenStudy (amistre64):

good, and the one for the 4?

OpenStudy (anonymous):

4x

OpenStudy (amistre64):

the usual notation for integration is; \[\int_{a}^{b}f(x)~dx=F(b)-F(a)\] you just found the F(x) parts soo \[\int_{2}^{3}8x^2+4~dx=(\frac83(3)^3+4(3))-(\frac83(2)^3+4(2))\]

OpenStudy (anonymous):

so the answer is 164/3

OpenStudy (amistre64):

thats what i get

OpenStudy (anonymous):

now one more question

OpenStudy (anonymous):

determine the region bounded by f(x)=\[4/\sqrt[2]{x}\] and the x axis between 16 and 25

OpenStudy (amistre64):

turn it into a power and follow the same rules you did for this one

OpenStudy (anonymous):

the anti derivitve of 4 will be 4x^2

OpenStudy (amistre64):

\[\frac{4}{\sqrt x}\] \[\frac{4}{x^{1/2}}\] \[f(x)=4x^{-1/2}\]

OpenStudy (anonymous):

ohh quotient rule?

OpenStudy (amistre64):

no, its just the reverse of the power rule, just like you did for that 8x^2

OpenStudy (amistre64):

add 1 and divide

OpenStudy (anonymous):

ooo

OpenStudy (amistre64):

\[\int kx^n=\frac {kx^{n+1}}{n+1} \] or recall\[\frac d{dx}(kx^n)=kn~x^{n-1}\] \[\int kn~x^{n-1}=\frac {kn~x^{n-1+1}}{n-1+1} =\frac {k\cancel{n}~x^{n}}{\cancel{n}}=kx^n\]

OpenStudy (anonymous):

answer =8

OpenStudy (amistre64):

8sqrt(25)-8sqrt(16) 8(5-4) ; i agree, its 8 :)